“Measurement of time intervals are affected by relative motion between an observer if you agree with the statement, explain the above statement.
1
Expert's answer
2020-10-12T05:27:00-0400

From the einstine's pastulate

time dilation "t=\frac{t_o}{\sqrt{1-\frac{v^2}{c^2}}}"

So, from this we can conclude that if the speed of the object which have the initial speed id v, now when it increase then value of the denominator will decrease and if the object's speed reach to the speed of the light then the time dilation will be infinite, so if speed of object will be small then the time dilation will be high.